Simplify:

√[144x10y12z18]
One last one, please show steps

Answer by sdmmadam@yahoo.com(530)

√[144x10y12z18]
=√{[(12)^2][(x^5)^2][(y^6)^2][(z^9)^2]} [using (a)^(mn) =(a^m)^n and here n=2 because we wish to find the sqroot and we know that {sqrt(t^2)}=t ----(*)]
= √{[(12)][(x^5)][(y^6)][(z^9)]}^2 (using {a^2b^2c^2c^2} = (abcd)^2)
={[(12)][(x^5)][(y^6)][(z^9)]} using (*)
=12x^5y^6z^9